OpenStudy (anonymous):

Please help! i only get decimals as the answer!.. The half-life of a certain radioactive material is 72 days. An initial amount of the material has a mass of 344 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth.

OpenStudy (amistre64):

... represents moving along into infinity, or the deletion of steps :/

OpenStudy (anonymous):

This is regular radioactive decay which is an exponential decay given by: \[N(t) = N _{0}e ^{-kt}\] Where you have some initial amount N0, a decay constant 'k', and some time exlapsed 't'

OpenStudy (amistre64):

A=Pe^(rt) ; divide of the P A/P = e^(rt) ; ln the sides ln(A/P) = rt ; divide off the t to find r ln(A/P) ------ = r ; since A = P/2 t ln(1/2) /t = r ; and t = 72 plug this back in as "rate" in: A = P e^(rt) to determine the rest is my thought

OpenStudy (amistre64):

i spose Amount and Principle dont apply in biology; but ... you know :)

OpenStudy (anonymous):

That's correct, only that since it's a decay, not a growth, the exponent is negative.

OpenStudy (amistre64):

ln(1/2) = -ln(2)

OpenStudy (anonymous):

@Lunarwolf58 It's really useful to remember that if T is the half-life (or the doubling time, for a growth equation), then kT = ln(2)